# NeuTSFlow: Modeling Continuous Functions Behind Time Series Forecasting

### v1

@@ 1 Introduction
Time series forecasting is a fundamental task in time series analysis, with critical applications spanning healthcare [8], climate science [24], and industrial systems [5]. Current forecasting methods [18, 30, 3] typically treat time series data as discrete sequences, focusing on learning mappings between historical and future observations [9, 26, 14, 2, 1, 16, 23]. This approach overlooks a key characteristic of real-world time series: they are discrete, noisy observations of underlying continuous processes, such as electricity load, wind speed, or temperature [25].
This naturally leads to the idea that if a model can learn the relationships between the functions underlying these discrete sequences, it would be able to capture the more fundamental characteristics of the underlying processes. Mathematically, however, discrete observations cannot uniquely determine an underlying continuous function. Even without noise, countless continuous functions can fit the same discrete points. When noise is present, this ambiguity increases, meaning a discrete time series represents not a single continuous function but a family of plausible functions consistent with the noisy observations.
Figure 1: Comparison of Traditional Predictive Models and the NeuTSFlow Prediction Paradigm
@@
To validate the effectiveness of our model and the function-family perspective, we conduct extensive evaluations on eight widely-used datasets across three tasks: Conventional forecasting task, Time Series Super-Resolution, and Cross-Resolution Temporal Learning. Time Series Super-Resolution evaluates the model’s ability to reconstruct high-resolution time series data from low-resolution inputs, while Cross-Resolution Temporal Learning assesses the model’s capability to leverage high-resolution data to assist in forecasting. The results underscore the effectiveness of both our model and the proposed methodology.
In summary, our key contributions are as follows:
@@
2 Preliminary
2.1 Reframing Time Series Forecasting and Relevant Notations
Let the time domain be a continuous interval $\mathcal{T}=[a,b]\subset\mathbb{R}$, where multivariate observations are obtained at discrete time points $\{t_{i}\}_{i=1}^{n}\subset\mathcal{T}$ with ordered timestamps $t_{1}<t_{2}<\cdots<t_{n}$. Each observation at time $t_{i}$ is a $C$-dimensional vector $\boldsymbol{y}_{i}\in\mathbb{R}^{C}$. Define historical sequence $\mathcal{X}=\{(t_{i_{j}},\boldsymbol{y}_{i_{j}})\}_{j=1}^{S}$, where $\{t_{i_{j}}\}$ is an ordered subsequence with $a\leq t_{i_{1}}<\cdots<t_{i_{s}}\leq b$. Forecast sequence $\mathcal{Y}=\{(t_{k_{q}},\boldsymbol{y}_{k_{q}})\}_{q=1}^{L}$, where $\{t_{k_{q}}\}$ satisfies $t_{i_{s}}<t_{k_{1}}<\cdots<t_{k_{l}}\leq b$. Traditional methods formulate the prediction task as learning a mapping
$G:\mathbb{R}^{S\times C}\to\mathbb{R}^{L\times C}.$ (1)
However, such predictive methods only learn the mapping between discrete sequences and fail to capture the continuous nature of the time series underlying the continuous processes. We now redefine time series prediction through the lens of continuous processes. Since finite discrete observations cannot uniquely determine a continuous function $u$, and given that time series constitute noisy observations of the underlying continuous process, a function space framework is introduced to effectively model process uncertainty. Define a real separable Hilbert space $\mathcal{F}$ of multivariate functions $u$, where each function $u\in\mathcal{F}$ is of the form: $u:\mathcal{T}\to\mathbb{R}^{c},\quad t_{k}\mapsto u(t_{k})=\big{(}u_{1}(t_{k}),\dots,u_{c}(t_{k})\big{)}^{\top}.$ To formalize the latent structure of continuous-time processes, we introduce a probability measure $\mu$ on the measurable space $(\mathcal{F},\mathcal{B}(\mathcal{F}))$, where $\mathcal{B}(\mathcal{F})$ denotes the Borel $\sigma$-algebra induced by the norm topology of $\mathcal{F}$. In the context of time-series prediction, we aim to model the transition between two probability measures: $\mu_{H}$, which encapsulates the statistical properties of historical time-series functions, and $\mu_{F}$, which represents the probability measure associated with predicted time-series functions. This transition captures the evolution of the underlying dynamics and uncertainty from the known historical data to the future, enabling us to make probabilistic predictions about the time - series values at the future time points.
$M:\mathcal{\mu}_{H}\to\mathcal{\mu}_{F}.$ (2)
2.2 Flow Matching
Let $X_{0}\sim p$ and $X_{1}\sim q$ be source and target distributions in $\mathbb{R}^{d}$. Flow Matching (FM) [9] constructs a probability path $p_{t}$ between $p_{0}=p$ and $p_{1}=q$ by learning a time-dependent velocity field $u_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ [26, 14, 2, 1, 16]. This field generates a flow $\psi_{t}$ satisfying the ODE: $\frac{d}{dt}\psi_{t}(x)=u_{t}(\psi_{t}(x)),\quad\psi_{0}(x)=x$ such that $X_{t}=\psi_{t}(X_{0})\sim p_{t}$. FM trains a neural network $u_{\theta}$ via regression loss:$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],X_{t}\sim p_{t}}\|u_{\theta}(t,X_{t})-u_{t}(X_{t})\|^{2}$.
@@ 2.2 Flow Matching
2.3 Neural Operator
Neural operators[12, 13, 20] learn mappings between infinite-dimensional function spaces, enabling data-driven solutions for PDEs and functional regression. Let $\mathcal{F}$ be a function space (e.g., $L^{2}(\Omega)$) and $\mu$ a probability measure over $\mathcal{F}$. A neural operator $G_{\theta}:\mathcal{F}\to\mathcal{F}$ approximates an operator $G^{\dagger}$ by minimizing:
@@ 3.1 Framework Construction
Figure 2: Training Process of NeuTSFlow
@@
3.1.1 Marginalization
While direct estimation of the velocity field is intractable by design, the marginal path can be explicitly constructed to derive its corresponding velocity field. We therefore perform the following marginalization[9]:
@@ 3.1.1 Marginalization
3.1.2 Path of Measure
Given the well-established absolute continuity of Gaussian measures in infinite-dimensional (separable) Banach spaces [21], we leverage this property to construct a conditional path of measures in this work, defined as:
@@ 3.1.2 Path of Measure
3.1.3 Prediction Target
Numerous studies [17, 28] in sequence prediction have investigated the reparameterization of prediction targets. Inspired by this, we design the output of the conditional velocity field to depend directly on the prediction target. As shown in Eq. 3, predicting $f$ is equivalent to predicting $v_{t}^{f,h}(g)$, since $h$ remains known during both training and inference. Thus, by defining $v_{t}^{f,h}(g)=f$, the velocity field effectively becomes a mapping that predicts the target sequence as a function $f$.
@@ 3.2 Marginal Velocity Field Model $u_{\theta}^{t}$
3.2.1 Velocity Field Model Training and Inference
Based on Eq. 3 and 4, our velocity field is conditioned on the functions at both endpoints. Therefore, during training, the input consists of the historical sequence’s corresponding function $h$, the function $g$ along the path at time $t$, and the time $t$ itself. The target sequence’s corresponding function $f$ serves as the training objective. Once this marginal velocity field is learned, inference becomes straightforward: given $h$, the marginal velocity field at any time $t$ can be computed. By solving the ODE $\frac{\partial}{\partial t}\phi_{t}(g)=v_{t}(\phi_{t}(g))$ using an ODE solver, the function $g$ at time $t$ can be determined. This process continues until $t=1$, at which point $f$ is obtained. The detailed training and inference procedure can be found in the Appendix A.3.
3.2.2 Model Inputs
Our Marginal Velocity Field Model is built upon the theory of neural operators [13, 12, 20], which ensures that we can learn the underlying relationships between functions through discrete input values. Let the discretized input associated with $g$ be $\mathcal{G}_{t}\in\mathbb{R}^{L\times C}$, and the discretized history sequence associated with $h$ be $\mathcal{H}\in\mathbb{R}^{S\times C}$, which serves as a conditional input to the model. The discretized predicted time series corresponding to $h$ is denoted as $\mathcal{Y}$.
@@ 3.2.2 Model Inputs
3.2.3 Normalization
To address the non-stationarity inherent in time series data, where the distribution evolves over time—a challenge not commonly faced in PDE-based solutions, we leverage the historical sequence to remove non-stationary components, specifically the mean and standard deviation. By incorporating these statistics as prior information for the prediction target, we enhance the model’s performance and robustness in dynamic environments [10].
@@ 3.2.3 Normalization
3.2.4 Spectral Decomposition
Time series data exhibit key characteristics such as long-term slow-varying trends, periodic patterns, and noise. To leverage these properties, we decompose the historical sequence $\mathcal{H}$ in the frequency domain into trend and seasonal components. The proposed framework is implemented using the one-dimensional Discrete Fourier Transform (DFT), applied to each input $\mathcal{H}$ as follows:
@@ 3.2.4 Spectral Decomposition
3.2.5 Dimension Embedding
We concatenate the input state $\mathcal{G}_{t}\in\mathbb{R}^{L\times C}$, history embedding $\mathcal{H}_{\text{emb}}\in\mathbb{R}^{L\times C}$, and temporal encoding $t\in\mathbb{R}$ along the channel dimension, constructing an augmented representation $\mathbf{z}_{0}=[\mathcal{G}_{t};\mathcal{H}_{\text{emb}};t]\in\mathbb{R}^{L\times(2C+1)}$. However, this embedding may insufficiently capture the mapping between original infinite-dimensional function spaces $\mathcal{F}\to\mathcal{G}$. We employ a dimension-expansion strategy through learnable linear projection: First, we append a new dimension to $\mathbf{z}_{0}$: $\mathbf{z}_{0,\text{expanded}}=\mathbf{z}_{0}\oplus\mathbf{1}_{L\times(2C+1)\times 1}\in\mathbb{R}^{L\times(2C+1)\times 1},$ where $\oplus$ denotes the addition of a singleton dimension. We then expand this dimension via a learnable projection matrix: $\mathbf{z}_{1}=\mathbf{z}_{0,\text{expanded}}\cdot W_{e}$, where $W_{e}\in\mathbb{R}^{1\times k},$ yielding $\mathbf{z}_{1}\in\mathbb{R}^{L\times(2C+1)\times k}$ with $k\gg 1$.
3.2.6 Spectral Temporal Learning
@@
Here, element-wise multiplication ($\odot$) models $\kappa(y,x)\cdot a(x)$ in frequency space, while summation over $i$ emulates the integral aggregation across input features. Retaining low-frequency modes ($m\leq M$) performs spectral filtering, analogous to kernel smoothing in the integral operator. The inverse RFFT (IRFFT) transforms the spectrum back to the temporal domain, **corresponding to the integral result $\int\kappa(y,x)a(x)\,\mathrm{d}x$: $\mathbf{Y_{\mathcal{R}}}=\text{IRFFT}(\mathbf{Y}_{\mathcal{F}},n=L)\in\mathbb{R}^{L\times(2C+1)\times k}.$ Finally, a channel-wise linear projection maps the $(2C+1)\times k$ features to $C$ dimensions: $\mathcal{Y}=\text{Linear}\left(\text{Flatten}(\mathbf{Y_{\mathcal{R}}},\text{start\_dim}=1)\right)\in\mathbb{R}^{L\times C}$ where the learnable projection weights aggregate information across channels and features to achieve the target dimensionality.
3.2.7 Function-to-Function Mapping Guarantee
Combining the aforementioned components, we arrive at the following proposition, which ensures the correctness of our neural operator theory and function space flow matching:
@@ 3.2.7 Function-to-Function Mapping Guarantee
The marginal velocity field model $u_{\theta}^{t}:\mathcal{F}\times[0,1]\to\mathcal{F}$ constitutes a neural operator that satisfies Function-to-Function Mapping. The operator $u_{\theta}^{t}$ is decomposed into the following interconnected modules: normalization, spectral decomposition, dimension embedding, and spectral temporal learning.
3.3 Loss Function
To perform functional regression on the marginal vector field, we aim to approximate $v_{t}(g)$ using a model $u_{t}^{\theta}(g)$ with parameters $\theta\in\mathbb{R}^{p}$. This can be achieved by minimizing the loss: $L(\theta)=\mathbb{E}_{t\sim U[0,1],g\sim\mu_{t}}\left[\|v_{t}(g)-u_{t}^{\theta}(g)\|^{2}\right],$ where $U[0,1]$ represents a uniform distribution over the interval $[0,1]$. The model $u$ is defined as a mapping $u:\mathbb{R}^{p}\times[0,1]\times F\to F$, i.e., a parameterized, time-dependent operator acting on the function space $F$. However, directly minimizing $L(\theta)$ is computationally intractable. Similar to [14], we consider the conditional loss:

### v2

@@ 1 Introduction
Time series forecasting is a fundamental task in time series analysis, with critical applications spanning healthcare [8], climate science [24], and industrial systems [5]. Current forecasting methods [18, 30, 3] typically treat time series data as discrete sequences, focusing on learning mappings between historical and future observations [9, 26, 14, 2, 1, 16, 23]. This approach overlooks a key characteristic of real-world time series: they are discrete, noisy observations of underlying continuous processes, such as electricity load, wind speed, or temperature [25]. This naturally leads to the idea that if a model can learn the relationships between the functions underlying these discrete sequences, it would be able to capture the more fundamental characteristics of the underlying processes. Mathematically, however, discrete observations cannot uniquely determine an underlying continuous function. Even without noise, countless continuous functions can fit the same discrete points. When noise is present, this ambiguity increases, meaning a discrete time series represents not a single continuous function but a family of plausible functions consistent with the noisy observations.
@@
Figure 1: Comparison between traditional time series forecasting and the NeuTSFlow paradigm. While traditional models learn mappings between discrete sequences, NeuTSFlow models the transition between the distributions of continuous functions corresponding to historical and future sequences, thereby capturing richer and more nuanced temporal dynamics.
@@
To validate the effectiveness of our model and the function-family perspective, we conduct extensive evaluations on eight widely-used datasets across three tasks: Conventional forecasting task, Time Series Super-Resolution, and Cross-Resolution Temporal Learning. Time Series Super-Resolution evaluates the model’s ability to reconstruct high-resolution time series data from low-resolution inputs, while Cross-Resolution Temporal Learning assesses the model’s capability to leverage high-resolution data to assist in forecasting. The results underscore the effectiveness of both our model and the proposed methodology. Our contributions are as follows.
@@
2 Preliminaries
To ground our proposed approach, we first revisit the foundational concepts behind time series forecasting. This includes a reframing of the prediction problem in function space, as well as key tools such as flow matching and neural operators.
Reframing Time Series Forecasting and Relevant Notations. Let the time domain be a continuous interval $\mathcal{T}=[a,b]\subset\mathbb{R}$, where multivariate observations are obtained at discrete time points $\{t_{i}\}_{i=1}^{n}\subset\mathcal{T}$ with ordered timestamps $t_{1}<t_{2}<\cdots<t_{n}$. Each observation at time $t_{i}$ is a $C$-dimensional vector $\boldsymbol{y}_{i}\in\mathbb{R}^{C}$. Define historical sequence $\mathcal{X}=\{(t_{i_{j}},\boldsymbol{y}_{i_{j}})\}_{j=1}^{S}$, where $\{t_{i_{j}}\}$ is an ordered subsequence with $a\leq t_{i_{1}}<\cdots<t_{i_{s}}\leq b$. Forecast sequence $\mathcal{Y}=\{(t_{k_{q}},\boldsymbol{y}_{k_{q}})\}_{q=1}^{L}$, where $\{t_{k_{q}}\}$ satisfies $t_{i_{s}}<t_{k_{1}}<\cdots<t_{k_{l}}\leq b$. Traditional methods formulate the prediction task as learning a mapping $G:\mathbb{R}^{S\times C}\to\mathbb{R}^{L\times C}.$ However, such predictive methods only learn the mapping between discrete sequences and fail to capture the continuous nature of the time series underlying the continuous processes. We now redefine time series prediction through the lens of continuous processes. Since finite discrete observations cannot uniquely determine a continuous function $u$, and given that time series constitute noisy observations of the underlying continuous process, a function space framework is introduced to effectively model process uncertainty. Define a real separable Hilbert space $\mathcal{F}$ of multivariate functions $u$, where each function $u\in\mathcal{F}$ is of the form: $u:\mathcal{T}\to\mathbb{R}^{c},\quad t_{k}\mapsto u(t_{k})=\big{(}u_{1}(t_{k}),\dots,u_{c}(t_{k})\big{)}^{\top}.$ To formalize the latent structure of continuous-time processes, we introduce a probability measure $\mu$ on the measurable space $(\mathcal{F},\mathcal{B}(\mathcal{F}))$, where $\mathcal{B}(\mathcal{F})$ denotes the Borel $\sigma$-algebra induced by the norm topology of $\mathcal{F}$. In the context of time-series prediction, we aim to model the transition between two probability measures: $\mu_{H}$, which encapsulates the statistical properties of historical time-series functions, and $\mu_{F}$, which represents the probability measure associated with predicted time-series functions. This transition captures the evolution of the underlying dynamics and uncertainty from the known historical data to the future, enabling us to make probabilistic predictions about the time - series values at the future time points. $M:\mathcal{\mu}_{H}\to\mathcal{\mu}_{F}.$
@@ 2 Preliminaries
Neural Operators. Neural operators[12, 13, 20] learn mappings between infinite-dimensional function spaces, enabling data-driven solutions for PDEs and functional regression. Let $\mathcal{F}$ be a function space (e.g., $L^{2}(\Omega)$) and $\mu$ a probability measure over $\mathcal{F}$. A neural operator $G_{\theta}:\mathcal{F}\to\mathcal{F}$ approximates an operator $G^{\dagger}$ by minimizing:
@@ 3.1 Framework Construction
Figure 2: Training overview of NeuTSFlow. The model learns a velocity field $u_{t}^{\theta}$ that aligns with the conditional velocity $v_{t}^{f,h}$ between sampled functions $h\sim\mu_{H}$ and $f\sim\mu_{F}$ along the path of measure $\mu_{t}=tf+(1-t)h$.
@@
Marginalization: Marginal Paths for Functional Flow Learning. Direct estimation of the velocity field is intractable by design. However, the marginal path can be constructed explicitly, enabling tractable approximation of the velocity field via functional integration. We therefore perform the following marginalization[9]:
@@
Measure-Path Construction for Infinite-Dimensional Flows. Given the well-established absolute continuity of Gaussian measures in infinite-dimensional (separable) Banach spaces [21], we leverage this property to construct a conditional path of measures in this work, defined as:
@@
Target Reparameterization for Functional Forecasting. Numerous studies [17, 28] in sequence prediction have investigated the reparameterization of prediction targets. Inspired by this, we design the output of the conditional velocity field to depend directly on the prediction target. As shown in (1), predicting $f$ is equivalent to predicting $v_{t}^{f,h}(g)$, since $h$ remains known during both training and inference. Thus, by defining $v_{t}^{f,h}(g)=f$, the velocity field effectively becomes a mapping that predicts the target sequence as a function $f$.
@@ 3.2 Marginal Velocity Field Model $u_{\theta}^{t}$
We now introduce a learnable model for the marginal velocity field $u_{\theta}^{t}$, which governs the transition between functions over time. This section details how the model is trained using discrete samples of functions and how it enables functional forecasting.
Velocity Field Model Training and Inference. Based on (1) and (2), our velocity field is conditioned on the functions at both endpoints. Therefore, during training, the input consists of the historical sequence’s corresponding function $h$, the function $g$ along the path at time $t$, and the time $t$ itself. The target sequence’s corresponding function $f$ serves as the training objective. Once this marginal velocity field is learned, inference becomes straightforward: given $h$, the marginal velocity field at any time $t$ can be computed. By solving the ODE $\frac{\partial}{\partial t}\phi_{t}(g)=v_{t}(\phi_{t}(g))$ using an ODE solver, the function $g$ at time $t$ can be determined. This process continues until $t=1$, at which point $f$ is obtained. The detailed training and inference procedure can be found in the Appendix A.3.
Model Inputs. Our Marginal Velocity Field Model is built upon the theory of neural operators [13, 12, 20], which ensures that we can learn the underlying relationships between functions through discrete input values. Let the discretized input associated with $g$ be $\mathcal{G}_{t}\in\mathbb{R}^{L\times C}$, and the discretized history sequence associated with $h$ be $\mathcal{H}\in\mathbb{R}^{S\times C}$, which serves as a conditional input to the model. The discretized predicted time series corresponding to $h$ is denoted as $\mathcal{Y}$.
@@
Normalization To address the non-stationarity inherent in time series data, where the distribution evolves over time—a challenge not commonly faced in PDE-based solutions, we leverage the historical sequence to remove non-stationary components, specifically the mean and standard deviation. By incorporating these statistics as prior information for the prediction target, we enhance the model’s performance and robustness in dynamic environments [10].
@@
Spectral Decomposition. Time series data exhibit key characteristics such as long-term slow-varying trends, periodic patterns, and noise. To leverage these properties, we decompose the historical sequence $\mathcal{H}$ in the frequency domain into trend and seasonal components. The proposed framework is implemented using the one-dimensional Discrete Fourier Transform (DFT), applied to each input $\mathcal{H}$ as follows:
@@
Dimension Embedding. We concatenate the input state $\mathcal{G}_{t}\in\mathbb{R}^{L\times C}$, history embedding $\mathcal{H}_{\text{emb}}\in\mathbb{R}^{L\times C}$, and temporal encoding $t\in\mathbb{R}$ along the channel dimension, constructing an augmented representation $\mathbf{z}_{0}=[\mathcal{G}_{t};\mathcal{H}_{\text{emb}};t]\in\mathbb{R}^{L\times(2C+1)}$. However, this embedding may insufficiently capture the mapping between original infinite-dimensional function spaces $\mathcal{F}\to\mathcal{G}$. We employ a dimension-expansion strategy through learnable linear projection: First, we append a new dimension to $\mathbf{z}_{0}$: $\mathbf{z}_{0,\text{expanded}}=\mathbf{z}_{0}\oplus\mathbf{1}_{L\times(2C+1)\times 1}\in\mathbb{R}^{L\times(2C+1)\times 1},$ where $\oplus$ denotes the addition of a singleton dimension. We then expand this dimension via a learnable projection matrix: $\mathbf{z}_{1}=\mathbf{z}_{0,\text{expanded}}\cdot W_{e}$, where $W_{e}\in\mathbb{R}^{1\times k},$ yielding $\mathbf{z}_{1}\in\mathbb{R}^{L\times(2C+1)\times k}$ with $k\gg 1$.
Spectral Temporal Learning.
@@
Here, element-wise multiplication ($\odot$) models $\kappa(y,x)\cdot a(x)$ in frequency space, while summation over $i$ emulates the integral aggregation across input features. Retaining low-frequency modes ($m\leq M$) performs spectral filtering, analogous to kernel smoothing in the integral operator. The inverse RFFT (IRFFT) transforms the spectrum back to the temporal domain, corresponding to the integral result $\int\kappa(y,x)a(x)\,\mathrm{d}x$: $\mathbf{Y_{\mathcal{R}}}=\text{IRFFT}(\mathbf{Y}_{\mathcal{F}},n=L)\in\mathbb{R}^{L\times(2C+1)\times k}.$ Finally, a channel-wise linear projection maps the $(2C+1)\times k$ features to $C$ dimensions: $\mathcal{Y}=\text{Linear}\left(\text{Flatten}(\mathbf{Y_{\mathcal{R}}},\text{start\_dim}=1)\right)\in\mathbb{R}^{L\times C}$ where the learnable projection weights aggregate information across channels and features to achieve the target dimensionality.
Function-to-Function Mapping Guarantee. Combining the aforementioned components, we arrive at the following proposition, which ensures the correctness of our neural operator theory and function space flow matching:
@@
The marginal velocity field model $u_{\theta}^{t}:\mathcal{F}\times[0,1]\to\mathcal{F}$ constitutes a neural operator that serves as the continuous-time version of DLinear [30] and satisfies Function-to-Function Mapping. The operator $u_{\theta}^{t}$ is decomposed into the following interconnected modules: normalization, spectral decomposition, dimension embedding, and spectral temporal learning.
Loss Function. To perform functional regression on the marginal vector field, we aim to approximate $v_{t}(g)$ using a model $u_{t}^{\theta}(g)$ with parameters $\theta\in\mathbb{R}^{p}$. This can be achieved by minimizing the loss: $L(\theta)=\mathbb{E}_{t\sim U[0,1],g\sim\mu_{t}}\left[\|v_{t}(g)-u_{t}^{\theta}(g)\|^{2}\right],$ where $U[0,1]$ represents a uniform distribution over the interval $[0,1]$. The model $u$ is defined as a mapping $u:\mathbb{R}^{p}\times[0,1]\times F\to F$, i.e., a parameterized, time-dependent operator acting on the function space $F$. However, directly minimizing $L(\theta)$ is computationally intractable. Similar to [14], we consider the conditional loss: